Using the z-table, find the corresponding area between 𝑧 = 0

and each of the following:

1. z = 0.92

2. z = 1.29

3. z = 2.73

4. z = βˆ’0.50

5. z = βˆ’2.98


1
Expert's answer
2021-03-31T13:34:17-0400

As we know from calculus, the area under the graph of a positive function "f(x)" between two values "x_1 <x_2" is equal to the integral "\int\limits_{x_1}^{x_2}f(x)dx =F(x_2)-F(x_1)" , where F(x) is any antidevative of the function "f(x)". If we take "f(x)=\frac{1}{\sqrt{2\pi}}e^{-x^2/2}" then its antiderivative is "\Phi(x)=\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^{x}e^{-t^2/2}dt", which is a cumulative function of distribution for standard normal distribution.


Using the table of CFD for standard normal distribution, we have

"\begin{matrix}
 z& -0.50 & 0.00 & 0.92& 1.29 & 2.73 & 2.98 \\
 \Phi(z)& 0.3085 & 0.5000 & 0.8212 & 0.9015 & 0.9968 & 0.9986
\end{matrix}"


The area under the graph of density of standard normal distribution between 𝑧 = 0 and "z=a" is:

if "a=-0.50" then "S=\Phi(0.00)-\Phi(-0.50)=0.5000-0.3085=0.1915"

if "a=0.92" then "S=\Phi(0.92)-\Phi(0.00)=0.8212-0.5000=0.3212"

if "a=1.29" then "S=\Phi(1.29)-\Phi(0.00)=0.9015-0.5000=0.4015"

if "a=2.73" then "S=\Phi(2.73)-\Phi(0.00)=0.9968-0.5000=0.4968"

if "a=2.98" then "S=\Phi(2.98)-\Phi(0.00)=0.9986-0.5000=0.4986"